Problem 3-10 A group of investors is considering buying the Wheelwright Corporation, but does not want to contribute to the companys financial support after the purchase. Wheelwrights management has offered the following financial statements covering last year (SM omitted): Wheelwright Corporation Balance Sheets ASSETS Beginning Ending Cash Accts receivable 13 20 CURRENT ASSETS 31 36 Fixed Assets Gross 100 115 (18) 97 133 Accumulated depreciation (1) Net fixed assets TOTAL ASSETS 119 LIABILITIES & EQUITY 21 Accts payable Accruals CURRENT LIABILITIES Debt 17 23 29 59Debt Equity TOTAL LIABILITIES & EQUITY 71 59 25 119 133 Wheelwright Corporation Income Statement Sales COGs Depreciation Gross Margin Expenses EBIT Interest EBT Tax Net income 98 35 57 20 37 28 20 Cost of Goods Sold Wheelwright paid no dividends and sold no new stock during the year. The firms tax rate is 40%. a. Develop Wheelwrights free cash flow and make a recommendation as to whether it seems to be an appropriate acquisition for the investors. Round your answer to one decimal place. Enter your answer in million of dollarInterest EBT Tax Net income 28 20 Cost of Goods Sold Wheelwright paid no dividends and sold no new stock during the year. The firms tax rate is 40%. a. Develop Wheelwrights free cash flow and make a recommendation as to whether it seems to be an appropriate acquisition for the investors. Round your answer to one decimal place. Enter your answer in million of dollar. b. Free Cash flow $ 15.4X millions The input in the box below will not be graded, but may be reviewed and considered by your instructor. C. Assume that the investors will purchase the company subject to its existing debt ($59M). Does that change your recommendation? No V Feedback ▼Check My Work Partially correct
Answer #1

Free cash flow = Net Income + interest expense - Tax shield on interest expense + Depreciation - Change in (Current assets - current liabilities) - Capital expenditure

Net Income = 20

Interest expense = 9

Tax shield on interest expense = Net interest expense * Tax rate = 9 *8/28 = 2.5714

Depreciation = 6

Current Assets - Current liabilities (beginning ) = 31-23 = 8

Current Assets - Current liabilities (end) = 36-29 = 7

Change in Working capital = 7-8 =-1

Capital expenditure = 97-88 =9

Free cash flow = 20+9 - 2.5714 -(-1) - 9 = 18.4

Free cash flow = 18.4